# Inference for multiple treatment effects using confounder importance learning

### v4

@@ 3.3. Confounder importance learning by Expectation-Propagation
In our examples ${\bm{\theta}}^{\textsc{eb}}$ and ${\bm{\theta}}^{\textsc{ep}}$ provided virtually indistinguishable inference, the latter incurring a significantly lower computational cost. Figure 9 shows a comparison for one of our simulated datasets. On the other hand, ${\bm{\theta}}^{\textsc{eb}}$ does provide slight advantages in some settings where the number of parameters $J+T$ was particularly large (see Section 4.1).
@@ 4. Results
In Section 4.1 we present a series of simulation studies, where we aim to illustrate the over-selection and under-selection issues discussed earlier across a range of settings. These range from no confounding settings, where all covariates are instruments, to full confounding scenarios where all covariates are confounders. We also consider single and multiple treatments, as well as varying sample sizes and problem dimensions. We next present two separate case studies. Section 4.2 studies the association between certain demographics and the hourly salary, and its evolution between 2010 and 2019 (prior to covid-19, to avoid potential pandemic-related distortions), to assess progress in wage discrimination. In Section 4.3 we analyse a putative association between less favorable environmental conditions at birth and subsequent crime levels some years later, following a study carried out by Donohue III and Levitt (2001), followed up by Belloni et al. (2014).
@@ 4.1.1. Single treatment
Figure 1 summarizes the results when the number of active covariates is $|{\bm{\gamma}}|_{0}=6$ out of a total of $J=49$, $n=100$, and the treatment effect is either strong ($\alpha=1$), weak ($\alpha=1/3$), or non-existent ($\alpha=0$). The two main features are as follows. First, standard high-dimensional methods such as LASSO and BMA incur a high RMSE in high-confounding settings, whereas methods such as DL and BAC that are designed to prevent omitted variable biases perform much better in this regime. Second, in low confounding settings DL and BAC incur a high RMSE, due to the over-selection of instruments into the outcome equation. ACPME mimics the same RMSE pattern as these two, overshooting the model size less than BAC, but similarly to DL (see Figure 11). In contrast, our CIL performs well at all levels of confounding. Note also that, when there truly is no treatment effect, CIL (and BMA, in some instances) attains a much lower RMSE than the oracle OLS. This occurs because CIL effectively shrinks the treatment estimate to zero. Of course, it is possible to modify methods such as BAC or ACPME to also run selection on the treatment and one would then expect a comparable shrinkage, our results simply point out the potential benefits in conducting selection on the treatment effects. The Empirical Bayes and the Expectation Propagation versions of CIL provide nearly indistinguishable results (not shown).
Figure 11 complements these results by showing the average size of the selected outcome model (top panels), and the probability of selecting the treatment (bottom panels). The results indicate that, as expected, the RMSE inflation incurred by LASSO and BMA in high-confounding settings is due to omitted variable biases.
@@
We next consider two extensions of our simulation study. First, Figure 12 considers a growing number of covariates, specifically $J+T=\{25,100,200\}$ with corresponding sample sizes $n=\{50,100,100\}$, in all cases under a strong treatment effect ($\alpha=1$). As dimensionality grows, the standard LASSO and BMA incur a significantly higher RMSE under strong confounding. Our CIL generally provides a significantly lower RMSE over BMA in high-confounding scenarios, and a similar RMSE under mild and no confounding. An exception is the larger $J+T=200$, where under mild and no confounding the RMSE for BMA was roughly half that for CIL, although the latter was still significantly better than DL and BAC. At $J+T=100$, ACPME departs from the behavior pattern of BAC and sensibly improves its relative performance for low levels of confounding, althought it cannot attain the results of CIL. At $J+T=200$, where $n<J+T$, ACPME cannot be computed. It is in this latter setting where we observe the only perceptible differences between the EB and EP approximations, with the former attaining better results, pointing to advantages of the EB approach in higher dimensions.
@@
As a second extension, Figure 13 shows the results when considering less sparse settings, specifically with $|{\bm{\gamma}}|_{0}=6$, 12 and 18 active parameters. Overall, the results are similar to Figures 1 and 12. Our CIL continued to provide a competitive and more robust behavior across levels of confounding, relative to the other considered methods. It is worth noting that again ACPME is able to improve the results of BAC, although it still cannot match the performance of CIL, particularly in low confounding scenarios.
@@ 4.1.2. Multiple treatments
To help understand under- and over-selection issues in multiple treatment inference, we consider an increasing number of treatments, with a maximum of $T=5$. There, every present treatment is active, setting $\alpha_{t}=1$ on all treatments. For all levels of $T$, we set $\beta_{j}=1$ for $j=1,\dots,20$, denoting the set of active covariates by ${\mathbf{x}}_{1:20}$, and $\beta_{j}=0$ for the rest of covariates ${\mathbf{x}}_{21:J}$. Regarding the association between treatments and covariates, ${\mathbf{x}}_{1:20}$ are divided into five disjoint subsets with four variables each, and each of these subsets is associated to a different treatment. The treatments depend linearly on the covariates of its associated subset. Additionally, each treatment also depends on a further subset of inactive covariates ${\mathbf{x}}_{21:J}$, i.e. instruments. In this case, the size of such subset is increasing by four with each added treatment: treatment 1 is associated to ${\mathbf{x}}_{21:24}$, treatment 2 is associated to ${\mathbf{x}}_{21:28}$, etc., up to treatment 5, which is associated to ${\mathbf{x}}_{21:40}$. All covariates affecting a treatment has a unit linear coefficient. The idea is that, as one considers a larger number of treatments $T$, one expects that potential ill-effects of over-selecting instruments and under-selecting confounders may become more problematic. Accordingly, as described our simulation considers $4T$ confounders and a growing number of instruments (4, 8, 12, 16, 20) for $T=(1,2,3,4,5)$. The rest of the design is as in Figure 1.
@@
where $\alpha_{t}^{*}$ is the data-generating truth and $\hat{\alpha}_{t}$ the estimate provided by each method. We observe similar trends as before. Methods prone to over-selection recover more instruments as $T$ increased. Some of these covariates are increasingly influential with $T$ as they are associated to more treatments, and so they become harder to discard. Interestingly, ACPME was designed to ameliorate the over-selection of instruments in multiple treatment settings, and indeed we observe an improved behavior as $T$ grows. Still, its RMSE ranged from 2 to 4 times larger than that of CIL.
@@ 4.2. Salary variation and discriminating factors
We analyze the USA Current Population Survey (CPS) microdata (Flood et al., 2020), which records many social, economic and job-related factors. We download data from 2010 and 2019 and analyze each year separately (see Section H.2 for details on data acquisition and pre-processing). We select individuals aged over 18, with a yearly income over \$1,000 and working 20-60 hours per week, giving $n=64,380$ and $n=58,885$ in 2010 and 2019, respectively. The covariates include characteristics of the place of residence, education, labor force status, migration status, household composition, housing type, health status, financial records, reception of subsidies, and sources of income (beyond wage). Overall, there are $J=278$ covariates, 228 given in the raw data plus 50 indicators for state. The outcome is the individual log-hourly wage, rescaled by the consumer price index of 1999, and we consider $T=4$ main treatments of interest: gender, black race, Hispanic ethnicity and Latin America as place of birth. These treatments are highly correlated to sociodemographic and job characteristics that can impact salary, i.e. there are many potential confounders. Since every state has its own regulatory, sociodemographic and political framework, we capture heterogeneous state effects by adding interactions for each pair of treatment and state. On these interactions, we apply a sum to zero constraint, so that the coefficients associated to the four treatments remain interpretable as average treatment effects across the USA, and the interactions as deviation from the average. Hence, overall we have $4+4\times 50=204$ parameters quantifying treatment effects, and our main interest is in the first four. To simplify computation in our CIL prior we assume a common $\theta_{t}$ shared between each main treatment and all its interactions with state, so that $\mbox{dim}({\bm{\theta}})=5$.
@@
Figure 4 reports the results for gender and race. More detailed results in Figure 10 show that none of the methods finds an association between salary variation and ethnicity or place of birth. The treatment effect for gender is picked up by all methods in both years with similar point estimates. All methods suggest a slight decrease of this effect in 2019. For race the methods vary in their findings. All methods find a negative association between black race and salary, but in 2019 OLS and DL estimate a fairly lower effect.
In order to understand this difference better, and explore whether it is due to over-selection, we analyze two additional augmented datasets where we add artificial instruments. Specifically, we incorporate 100 instruments in the first scenario, and 200 in the second one. The instruments are split into four equal subsets, each of which is designed to correlate to one of the four main treatments, see Section H.3 for full details. The resulting average correlation between gender and its associated artificial instruments is $0.83$, and analogously $0.69$, $0.76$ and $0.67$ for black race, ethnicity and place of birth. Upon adding said instruments, the confidence intervals for OLS and DL become notably wider, whereas CIL and BMA results remain particularly robust. This variance inflation is particularly pronounced for the effect of black race, which in 2019 lead to a loss of statistical significance according to OLS and DL. These findings suggest that the smaller racial gap estimated in 2019 in the original data may be due to variance inflation rather than an actual improvement in the racial gap.
The full scope of a Bayesian inferential framework is materialized when, additionally to quantifying treatment effects, one also considers more complex functions of the parameters. As an illustration, we study a measure of overall treatment contribution to deviations from the average salary. The idea is that the conditional associations between salary and the four treatments (gender, race, Hispanic ethnicity and birth in Latin America) may reflect salary discrimination associated to these demographics, and it is hence interesting to quantify the overall effect of all four treatments. For a new observation $n+1$, with observed treatments $\mathbf{d}_{n+1}$ and covariates $\mathbf{x}_{n+1}$, let

### v6

@@ 3.3. Confounder importance learning by expectation-propagation
In most of our examples ${\bm{\theta}}^{\textsc{eb}}$ and ${\bm{\theta}}^{\textsc{ep}}$ provided virtually indistinguishable inference, the latter incurring a significantly lower computational cost. Figure S4 shows a comparison for one of our simulated datasets. On the other hand, ${\bm{\theta}}^{\textsc{eb}}$ does provide slight advantages in some settings where the number of parameters $J+T$ was particularly large (see Section 5.1).
4. Asymptotic analysis
Empirical Bayes seeks ${\bm{\theta}}^{\textsc{eb}}$ such that the prior mean of the confounding coefficient $E(\kappa_{t}\mid{\bm{\theta}}^{\textsc{eb}})$ equals its posterior mean $E(\kappa_{t}\mid{\mathbf{y}},{\bm{\theta}}^{\textsc{eb}})$ (see Section 3.2), and a similar interpretation applies to the expectation-propagation estimator ${\bm{\theta}}^{\textsc{ep}}$. Theorem 4.1 shows that $E(\kappa_{t}\mid{\mathbf{y}},{\bm{\theta}})$ converges in probability to the true $\kappa_{t}^{*}$ uniformly across ${\bm{\theta}}$, as $n\rightarrow\infty$, which implies that $E(\kappa_{t}\mid{\bm{\theta}}^{EB})\stackrel{{\scriptstyle P}}{{\longrightarrow}}\kappa_{t}^{*}$. That is, that empirical Bayes sets the CIL prior to match the true confounding coefficient, asymptotically. We allow for model (1) to be misspecified, and focus on the case where $p$ is fixed. High-dimensional settings where $p\gg n$ are also interesting, but they require a delicate treatment beyond our scope.
From (12), to prove that $E(\kappa_{t}\mid{\mathbf{y}},{\bm{\theta}})\stackrel{{\scriptstyle P}}{{\longrightarrow}}\kappa_{t}^{*}$ it suffices that ${\bf f}_{t}\stackrel{{\scriptstyle P}}{{\longrightarrow}}{\bf f}_{t}^{*}$ and that $p({\bm{\gamma}}\mid{\mathbf{y}},{\bm{\theta}})$ converges to a point mass distribution at ${\bm{\gamma}}^{*}$. The convergence of ${\bf f}_{t}$ follows from standard theory. For example, if ${\bf f}_{t}$ is obtained from the MLE separately regressing each treatment on the covariates, then the result holds under mild conditions, even if the model for the treatments is misspecified (van der Vaart (1998), Theorem 5.7). If ${\bf f}_{t}$ is obtained from LASSO regressions, then consistency also holds under mild conditions, see Bühlmann and van de Geer (2011) (Chapters 6-7). Regarding $p({\bm{\gamma}}\mid{\mathbf{y}},{\bm{\theta}})$, its convergence to ${\bm{\gamma}}^{*}$ can be established by showing that the posterior odds between any model ${\bm{\gamma}}$ and ${\bm{\gamma}}^{*}$ converges to 0. This requires mild conditions D0-D3, see Section S7.1. These conditions require log-likelihood concavity at the MLE (which holds for full-rank generalized GLMs under the canonical link), that the MLE is consistent as $n\rightarrow\infty$, that the asymptotic hessian is strictly positive definite, and a betamin condition that can be relaxed but simplifies our exposition.
Theorem 4.1.
Suppose that Conditions D0-D3 hold and that ${\bf f}_{t}\stackrel{{\scriptstyle P}}{{\longrightarrow}}{\bf f}_{t}^{*}$ as $n\rightarrow\infty$. Then, $\sup_{{\bm{\theta}}}|E(\kappa_{t}\mid{\mathbf{y}},{\bm{\theta}})-\kappa_{t}^{*}|\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$, as $n\rightarrow\infty$.
We next sketch the proof, see Section S7 for further details. We consider separately overfitted and non-overfitted models. The former refer to models $({\bm{\delta}},{\bm{\gamma}})$ that include all parameters in $({\bm{\delta}}^{*},{\bm{\gamma}}^{*})$ plus some truly zero parameters. In contrast, non-overfitted models fail to include some truly non-zero parameters.
We denote by $d=|({\bm{\delta}},{\bm{\gamma}})|_{0}-|({\bm{\delta}}^{*},{\bm{\gamma}}^{*})|_{0}$ the difference between model dimensions, by $d_{1}=\sum_{j=1}^{J}\gamma_{j}(1-\gamma_{j}^{*})$ the number of covariates included in ${\bm{\gamma}}$ but not in ${\bm{\gamma}}^{*}$, and by $d_{2}=\sum_{j=1}^{J}(1-\gamma_{j})\gamma_{j}^{*}$ that of covariates included in ${\bm{\gamma}}^{*}$ but not in ${\bm{\gamma}}$. Section S7 shows that, for overfitted models, the posterior odds satisfy
(17) $\frac{p({\bm{\delta}},{\bm{\gamma}}\mid{\mathbf{y}},{\bm{\theta}})}{p({\bm{\delta}}^{*},{\bm{\gamma}}^{*}\mid{\mathbf{y}},{\bm{\theta}})}\leq\left(\frac{\bar{\rho}}{(n\tau)^{3/2}(1-\bar{\rho})}\right)^{d}\times O_{p}(1),$
as $n\rightarrow\infty$, uniformly across ${\bm{\theta}}$. This polynomial rate in $n$ reflects the effect of using a pMOM prior, for any other prior with a density that does not vanish at zero the rate is slower, specifically $(n\tau)^{3/2}$ is replaced by $(n\tau)^{1/2}$. Note that for our default upper-bound on the prior inclusion probability $\bar{\rho}=0.95$, (17) converges to 0, as desired.
In contrast, for non-overfitted models,
$\log\left(\frac{p({\bm{\delta}},{\bm{\gamma}}\mid{\mathbf{y}},{\bm{\theta}})}{p({\bm{\delta}}^{*},{\bm{\gamma}}^{*}\mid{\mathbf{y}},{\bm{\theta}})}\right)\leq-\frac{3d}{2}\log(n\tau)-nc+d_{1}\log\left(\frac{\bar{\rho}}{1-\bar{\rho}}\right)+d_{2}\log\left(\frac{1-\underline{\rho}}{\underline{\rho}}\right)+O_{p}(1)$
again uniformly in ${\bm{\theta}}$. That is, for overfitted models one obtains a faster rate that is almost exponential in $n$. For these posterior odds to vanish it suffices that
$\lim_{n\rightarrow\infty}-\frac{3(|{\bm{\delta}}^{*}|_{0}+|{\bm{\gamma}}^{*}|_{0})}{2}\log(n\tau)+nc-d_{1}\log\left(\frac{\bar{\rho}}{1-\bar{\rho}}\right)+|{\bm{\gamma}}^{*}|_{0}\log(\underline{\rho})=\infty,$
where we used that $d\geq-(|{\bm{\delta}}^{*}|_{0}+|{\bm{\gamma}}^{*}|_{0})$ and $d_{2}\leq|{\bm{\gamma}}^{*}|_{0}$, and recall that our default is $\underline{\rho}=1/J$. Although we do not consider high-dimensional theory here, note that there one often sets sparse prior probabilities $\bar{\rho}<1/2$, then $-d_{1}\log(\bar{\rho}/[1-\bar{\rho}])>0$, which helps ensure that the sufficient condition above is met.
@@ 5. Results
In Section 5.1 we present a series of simulation studies, where we aim to illustrate the over-selection and under-selection issues discussed earlier across a range of settings. These range from no confounding settings, where all covariates are instruments, to full confounding scenarios where all covariates are confounders. We also consider single and multiple treatments, as well as varying sample sizes and problem dimensions. We next present two separate case studies. Section 5.2 studies the association between certain demographics and the hourly salary, and its evolution between 2010 and 2019 (prior to covid-19, to avoid potential pandemic-related distortions), to assess wage discrimination. In Section 5.3 we analyse a putative association between less favorable environmental conditions at birth and subsequent crime levels some years later, following a study carried out by Donohue III and Levitt (2001), retaken by Belloni et al. (2014).
@@ 5.1.1. Single treatment
Figure 1 summarizes the results when the number of active covariates is $|{\bm{\gamma}}|_{0}=6$ out of a total of $J=49$, $n=100$, and the treatment effect is either strong ($\alpha=1$), weak ($\alpha=1/3$), or non-existent ($\alpha=0$). The two main features are as follows. First, standard high-dimensional methods such as LASSO and BMA incur a high RMSE in high-confounding settings, incurring both high bias and variance (Figure S6)whereas methods such as DL and BAC that are designed to prevent omitted variable biases perform much better in this regime. Second, in low confounding settings DL, BAC and ACPME incur a high RMSE, due to high variance (Figure S6). Figure S10 shows that this is due to selecting all instruments, resulting in a larger model size (Figure S11). Figure S7 shows that for BAC this behavior is highly sensitive to the choice of hyper-parameter $\omega$, driving the prior dependence between inclusion in the outcome and treatment equations ($\omega=\infty$ for complete dependence, $\omega=1$ for independence). In contrast, our CIL framework performs well at all levels of confounding, by avoiding the inclusion of instruments. Accordingly, in low-confounding scenarios we obtain hyper-parameter estimates $\hat{\theta}_{1}<0$, and $\hat{\theta}_{1}>0$ under high-confounding (Figure S8). It is also informative to assess how the prior inclusion odds assigned by CIL behave relative to those of ACPME. Figure S9 shows that in high confounding scenarios CIL assigns higher prior inclusion odds to controls that are associated to the treatment (which are mostly confounders) than it does in low confounding (when they are mostly instruments). In contrast, ACPME sets the same prior odds in either high or low confounding, i.e. it does not adapt to the true amount of confounding in the data. Also, the CIL prior inclusion odds are overall smaller, this is because $\theta_{0}$ in (4) learns the true amount of sparsity, as shown by our asymptotic study in Section 4. For further discussion see Section S8.6.
@@
We remark that, when there truly is no treatment effect, CIL (and BMA, in some instances) attains a much lower RMSE than the oracle OLS. This occurs because CIL effectively shrinks the treatment estimate to zero. Of course, it is possible to modify methods such as BAC or ACPME to also run selection on the treatment and one would then expect a comparable shrinkage, our results simply point out the potential benefits in conducting selection on the treatment effects. The Empirical Bayes and the expectation-propagation versions of CIL provide nearly indistinguishable results (not shown). Figure S11 complements these results by showing the probability of selecting the treatment (bottom panels). Overall, the RMSE inflation incurred by LASSO and BMA in high-confounding settings is due to omitted variable biases, and that of double LASSO, BAC AND ACPME under low confounding is due to selecting instruments.
@@
We next consider two extensions of our simulation study. First, Figure S12 considers a growing number of covariates, specifically $J+T=\{25,100,200\}$ with corresponding sample sizes $n=\{50,100,100\}$, in all cases under a strong treatment effect ($\alpha=1$). As dimensionality grows, the standard LASSO and BMA incur a significantly higher RMSE under strong confounding. Our CIL generally provides a significantly lower RMSE over BMA in high-confounding scenarios, and a similar RMSE under mild and no confounding. An exception is the larger $J+T=200$, where under mild and no confounding the RMSE for BMA was roughly half that for CIL, although the latter was still significantly better than DL and BAC. At $J+T=100$, ACPME departs from the behavior pattern of BAC and sensibly improves its relative performance for low levels of confounding, although it cannot attain the results of CIL. At $J+T=200$, where $n<J+T$, ACPME cannot be computed. It is in this latter setting where we observe the only perceptible differences between the EB and EP approximations, with the former attaining better results, pointing to advantages of the EB approach in higher dimensions.
@@
As a second extension, Figure S13 shows the results when considering less sparse settings, specifically with $|{\bm{\gamma}}|_{0}=6$, 12 and 18 active parameters. Overall, the results are similar to Figures 1 and S12. Our CIL continues to provide a competitive and more robust behavior across levels of confounding, relative to the other considered methods. It is worth noting that again ACPME is able to improve the results of BAC, although it still cannot match the performance of CIL, particularly in low confounding scenarios.
@@ 5.1.2. Multiple treatments
To help understand under- and over-selection issues in multiple treatment inference, we consider an increasing number of treatments, with a maximum of $T=5$. There, every present treatment is active, setting $\alpha_{t}=1$ on all treatments. For all levels of $T$, we set $\beta_{j}=1$ for $j=1,\dots,20$, denoting the set of active covariates by ${\mathbf{x}}_{1:20}$, and $\beta_{j}=0$ for the rest of covariates ${\mathbf{x}}_{21:J}$. Regarding the association between treatments and covariates, ${\mathbf{x}}_{1:20}$ are divided into five disjoint subsets with four variables each, and each of these subsets is associated to a different treatment. The treatments depend linearly on the covariates of its associated subset. Additionally, each treatment also depends on a further subset of inactive covariates ${\mathbf{x}}_{21:J}$, i.e. instruments. In this case, the size of such subset is increasing by four with each added treatment: treatment 1 is associated to ${\mathbf{x}}_{21:24}$, treatment 2 is associated to ${\mathbf{x}}_{21:28}$, etc., up to treatment 5, which is associated to ${\mathbf{x}}_{21:40}$. All covariates that affect a treatment have a regression coefficient equal to 1. The idea is that, as one considers a larger number of treatments $T$, one expects that potential ill-effects of over-selecting instruments and under-selecting confounders may become more problematic. Accordingly, as described our simulation considers $4T$ confounders and a growing number of instruments (4, 8, 12, 16, 20) for $T=(1,2,3,4,5)$. The rest of the design is as in Figure 1.
@@
where $\alpha_{t}^{*}$ is the data-generating truth and $\hat{\alpha}_{t}$ the estimate provided by each method. We observe similar trends as before. Methods prone to over-selection recover more instruments as $T$ increases. Some of these covariates are increasingly influential with $T$ as they are associated to more treatments, and so they become harder to discard. Interestingly, ACPME was designed to ameliorate the over-selection of instruments in multiple treatment settings, and indeed we observe an improved behavior as $T$ grows. Still, its RMSE ranged from 2 to 4 times larger than that of CIL.
@@ 5.2. Salary variation and discriminating factors
We analyze the USA Current Population Survey (CPS) microdata (Flood et al., 2020), which records many social, economic and job-related factors. We download data from 2010 and 2019 and analyze each year separately (see Section S8.2 for details on data acquisition and pre-processing). We select individuals aged over 18, with a yearly income over \$1,000 and working 20-60 hours per week, giving $n=64,380$ and $n=58,885$ in 2010 and 2019, respectively. The covariates include characteristics of the place of residence, education, labor force status, migration status, household composition, housing type, health status, financial records, reception of subsidies, and sources of income (beyond wage). Overall, there are $J=278$ covariates, 228 given in the raw data plus 50 indicators for state. The outcome is the individual log-hourly wage, rescaled by the consumer price index of 1999, and we consider $T=4$ main treatments of interest: sex, black race, Hispanic ethnicity and Latin America as place of birth. Specifically, we introduce a female indicator for individuals who declared female as their single sex, and a black indicator for those who declared black as their single race. These treatments are highly correlated to sociodemographic and job characteristics that can impact salary, i.e. there are many potential confounders. Since every state has its own regulatory, sociodemographic and political framework, we capture heterogeneous state effects by adding interactions for each pair of treatment and state. On these interactions, we apply a sum to zero constraint, so that the coefficients associated to the four treatments remain interpretable as average treatment effects across the USA, and the interactions as deviation from the average. Hence, overall we have $4+4\times 50=204$ parameters quantifying treatment effects, and our main interest is in the first four. To simplify computation in our CIL prior we assume a common $\theta_{t}$ shared between each main treatment and all its interactions with state, so that $\mbox{dim}({\bm{\theta}})=5$. Since there are 4 treatments and their $4\times 50=200$ interactions with states, in principle $\mbox{dim}({\bm{\theta}})=205$ (including $\theta_{0}$). We view this as undesirable because optimizing over a 205-dimensional state would create a significant bottleneck, and because one does not expect to estimate precisely so many parameters. Further, one expects some structure in ${\bm{\theta}}$. If there is high confounding for a treatment in one state (positive entry in ${\bm{\theta}}$ for that state), then the same may hold in many other states. Setting $\mbox{dim}({\bm{\theta}})=5$ assumes that such measure of confounding is the same across all states.
@@
Figure 4 reports the results for sex and race. More detailed results in Figure S5 show that none of the methods finds an association between salary variation and ethnicity or place of birth. The treatment effect for sex is picked up by all methods in both years with similar point estimates. All methods suggest a slight decrease of this effect in 2019. For race the methods vary in their findings. All methods find a negative association between black race and salary, but in 2019 OLS and DL estimate a fairly lower effect.
In order to understand this difference better, and explore whether it is due to over-selection, we analyze two additional augmented datasets where we add artificial instruments. Specifically, we incorporate 100 instruments in the first scenario, and 200 in the second one. The instruments are split into four equal subsets, each of which is designed to correlate to one of the four main treatments, see Section S8.3 for full details. The resulting average correlation between sex and its associated artificial instruments is $0.83$, and analogously $0.69$, $0.76$ and $0.67$ for black race, ethnicity and place of birth. Upon adding said instruments, the confidence intervals for OLS and DL become notably wider, whereas CIL and BMA results remain particularly robust. This variance inflation is particularly pronounced for the effect of black race, which in 2019 lead to a loss of statistical significance according to OLS and DL. These findings suggest that the smaller racial gap estimated in 2019 in the original data may be due to variance inflation rather than an actual improvement in the racial gap.
The full scope of a Bayesian inferential framework is materialized when, additionally to quantifying treatment effects, one also considers more complex functions of the parameters. As an illustration, we study a measure of overall treatment contribution to deviations from the average salary. The idea is that the conditional associations between salary and the four treatments ( sex , race, Hispanic ethnicity and birth in Latin America) may reflect salary discrimination associated to these demographics, and it is hence interesting to quantify the overall effect of all four treatments. For a new observation $n+1$, with observed treatments $\mathbf{d}_{n+1}$ and covariates $\mathbf{x}_{n+1}$, let